OpenStudy (anonymous):

Write the equation of the line that passes through (-1, 4) with a slope of -1 in standard form. -x + 4y = -1 x - 4y = -1 x + y = 3 x - y = -3 Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. 4x + 4y = 6 x + 4y = 12 4x + 4y = -12 x - 4y = -6 Write the equation of the line that passes through (3, 6) and (0, -3) in slope-intercept form. y = 3x - 6 y = 3x - 3 y = 3x + 3 y = 3x + 6

OpenStudy (nurali):

First y-y1=m(x-x1) m=slope (x,y1)=(-1,4)

OpenStudy (nurali):

put value in formula and tell me what answer.

OpenStudy (nurali):

(x1,y1)=(-1,4)

OpenStudy (anonymous):

i dont kno thats y im asking! lol

OpenStudy (nurali):

y-(4)=-1(x-(-1)) y-4=-x-1 y=-x+3 x+y=3

OpenStudy (nurali):

Second Let (2, 2) and (6, 3) be (x1,y1) & (x2,y2) First find the slope m=(y2-y1)/(x2-x1) m=1/4 Then take any pt (2, 2) or (6, 3) (y-y1)=m(x-x1) Then substitute the values of m & x1&y1 x-4y+6=0

OpenStudy (anonymous):

can u jus give me the numbers and letters??

OpenStudy (anonymous):

well the numbers n answers

OpenStudy (anonymous):